# Tensor-polarized twist-3 parton distribution functions $f_{LT}(x)$ for the spin-1 deuteron by using twist-2 relations

### v1

@@ 1 Introduction
The JLab experiments for the tensor-polarized deuteron will be done in the near future. $Q^{2}(=-q^{2})$ values, where $q$ is the momentum transfer in the electron scattering, are not much larger than the typical hadronic energy scale of the order of 1 GeV2 in the JLab measurements. This fact suggests that higher-twist terms should be understood in extracting even the twist-2 tensor-polarized structure functions. Now, the theoretical formalism is completed to calculate cross sections of semi-inclusive DIS (SIDIS) processes for spin-1 hadrons [25], and these SIDIS processes could be used for measuring the tensor-polarized twist-3 distribution functions $f_{LT}(x)$. In addition, Drell-Yan processes could be used for studying $f_{LT}(x)$ at the hadron accelerator facilities [26]. Considering this situation, we show possible twist-3 distribution functions $f_{LT}(x)$ theoretically for the deuteron in this work. Because the distributions $f_{LT}(x)$ are written in terms of the twist-2 distributions $f_{1LL}(x)$ by the WW-like relation and a parametrization exists for $f_{1LL}(x)$, actually tensor-polarized PDFs $\delta_{T}q(x)$ [27], we can calculate $f_{LT}(x)$ by using this twist-2 relation. The obtained $f_{LT}(x)$ should satisfy the BC-like sum rule.
@@ 2 Twist-2 and 3 tensor-polarized parton distribution functions of the spin-1 deuteron
Here, $k$, $P$, and $T$ are the quark momentum, the deuteron momentum, and the tensor polarization of the deuteron, $\psi$ is the quark field, and $W^{[c]}$ is the gauge link with the integral path $c$ necessary for the color gauge invariance. Then, integrating the function over the momenta $k^{-}$ and the transverse momentum $\vec{k}_{T}$ and fixing the $k^{+}$ component, we obtain the collinear correlation function
@@
$\Phi(x_{D})=\frac{1}{2}\,\bigg{[}$ $\,S_{LL}\,\not{\bar{n}}\,f_{1LL}(x_{D})+\frac{M_{D}}{P^{+}}\,S_{LL}\,e_{LL}(x_{D})$
@@
[…]
The polarization parameters for the spin-1 deuteron are given in the spin vector $\vec{S}$ and tensor $T_{ij}$ defined by the polarization vector $\vec{E}$ of the spin-1 hadron as [18, 19]
$\vec{S}$ $=\text{Im}\,(\,\vec{E}^{\,*}\times\vec{E}\,)=(S_{T}^{x},\,S_{T}^{y},\,S_{L}),$
$T_{ij}$ $=\frac{1}{3}\delta_{ij}-\text{Re}\,(\,E_{i}^{\,*}E_{j}\,)$
$=\frac{1}{2}\begin{pmatrix}-\frac{2}{3}S_{LL}+S_{TT}^{xx}&S_{TT}^{xy}&S_{LT}^{x}\\ S_{TT}^{xy}&-\frac{2}{3}S_{LL}-S_{TT}^{xx}&S_{LT}^{y}\\ S_{LT}^{x}&S_{LT}^{y}&\frac{4}{3}S_{LL}\end{pmatrix}$ (5)
Then, a twist-3 matrix element is expressed by the tensor-polarized PDFs as [20]
$\xi_{\mu}\langle\,P,T\left|\,\bar{\psi}(0)(\partial^{\mu}\gamma^{\alpha}-\partial^{\alpha}\gamma^{\mu})\psi(\xi)\,\right|P,T\,\rangle$
$=2M_{D}S_{LT}^{\,\alpha}\int_{-1}^{1}dx_{D}\,e^{-ix_{D}P^{+}\xi^{-}}\left[-\frac{3}{2}\,f_{1LL}(x_{D})+f_{LT}(x_{D})\right$.
$\left.\hskip 122.34692pt-\frac{d}{dx_{D}}x_{D}f_{LT}(x_{D})\right],$ (6)
where the negative-$x_{D}$ distributions are related to the antiquark ones [20].
On the other hand, this matrix element is also expressed by the field tensor $G^{\mu\nu}$ and its dual one $\tilde{G}^{\mu\nu}$ as
$\xi_{\mu}\,\bar{\psi}(0)\big{(}\partial^{\mu}\gamma^{\alpha}-\partial^{\alpha}\gamma^{\mu}\big{)}\psi(\xi)\\ =g\int^{1}_{0}dt\,\bar{\psi}(0)\!\left[\,i\left(t-\frac{1}{2}\right)G^{\alpha\mu}(t\xi)-\frac{1}{2}\gamma_{5}\widetilde{G}^{\alpha\mu}(t\xi)\,\right]\xi_{\mu}\not{\xi}\,\psi(\xi),$ (7)
where the variable $t$ is introduced for expressing the gluon field by the field strength tensor [20]. The matrix element of the right-hand side of this equation is related to the multiparton distribution functions $F_{G,LT}$ and $G_{G,LT}$ as
$\int\frac{d(P\cdot\xi)}{2\pi}\,e^{ix_{D1}P\cdot\xi}\,\bigg{\langle}\,P,T\,\bigg{|}\,g\int^{1}_{0}dt\,\bar{\psi}(0)\left[i\left(t-\frac{1}{2}\right)G^{\alpha\mu}(t\xi)\right$.
$\left.\hskip 105.2751pt-\frac{1}{2}\gamma_{5}\,\tilde{G}^{\alpha\mu}(t\xi)\right]\xi_{\mu}\not{\xi}\psi(\xi)\,\bigg{|}\,P,T\,\bigg{\rangle}_{\xi^{+}=\vec{\xi}_{T}=0}$
$=-2M_{D}S_{LT}^{\alpha}{\cal P}\!\int_{-1}^{1}\!dx_{D2}\frac{1}{x_{D1}-x_{D2}}$
$\hskip 39.83368pt\times\left\{\frac{\partial}{\partial x_{D1}}\left[F_{G,LT}(x_{D1},x_{D2})+G_{G,LT}(x_{D1},x_{D2})\right]\right$.
$\left.\hskip 48.36958pt+\frac{\partial}{\partial x_{D2}}\left[F_{G,LT}(x_{D2},x_{D1})+G_{G,LT}(x_{D2},x_{D1})\right]\right\},$ (8)
where $\cal P$} indicates the principal value integral. Now, the twist-3 effects are expressed by the multiparton distribution functions. If these twist-3 effects are neglected, Eqs. (2), (7), and (2) lead to the WW-like relation
@@
and the BC-like sum rule
$\int_{0}^{1}dx_{D}\,f_{2LT}^{\,q+}(x_{D})=0,\ \ \ f_{2LT}^{\,q+}(x)\equiv\frac{2}{3}f_{LT}^{\,q+}(x_{D})-f_{1LL}^{\,q+}(x_{D}),$ (10)
for the spin-1 PDFs. For a review of the original WW relation and the BC sum rule in the spin-1/2 case, one may look at Ref. [28]. Here, the quark flavor type is shown by $q$ in the superscript as $f^{\,q+}$, which is defined by
@@ 3.2 Twist-3 distribution functions $f_{LT}(x)$ by using the twist-2 relations
Using the PDFs in Fig. 1, we calculated the tensor-polarized twist-2 distribution functions $f_{1LL}^{\,q+}$, which are expressed by the PDFs $\delta_{T}q$ and $\delta_{T}\bar{q}$ as [1, 20]
@@
In order to confirm the BC-like sum rule in Eq. (20), the distributions $f_{1LL}^{\,q/D+}(x)$ and $f_{LT}^{\,q/D+}(x)$ are combined to obtain to $f_{2LT}^{\,q+}(x)$ at $Q^{2}=2.5$ GeV2 in Fig. 4. According to the sum rule, the integrals of these distributions should vanish. In fact, the positive and negative areas in this figure cancel with each other. However, finite distributions still exist at $x\leq 10^{-3}$ in Fig. 4 and the integrals become much smaller, of the order of $10^{-5}$, if the integrals are calculated from $x_{min}=10^{-9}$.
The higher-twist distributions are generally suppressed by $1/Q$ in cross sections in comparison with the leading-twist ones. However, the $Q^{2}$ values of the JLab experiment are not very large, so that there is a possibility that the higher-twist distributions $f_{LT}^{\,q/D+}(x)$ could be measured experimentally by the semi-inclusive DIS [25] in addition to the leading-twist ones $f_{1LL}^{\,q/D+}(x)$ [3]. Furthermore, Drell-Yan processes could be used for finding the twist-3 distributions $f_{LT}^{\,q/D+}(x)$ [26].
@@ 4 Summary
Using the parametrization for the tensor-polarized PDFs $\delta_{T}q^{D}(x)$ and $\delta_{T}\bar{q}^{D}(x)$, we calculated the tensor-polarized twist-3 PDFs $f_{LT}^{\,u/D+}(x)=f_{LT}^{\,d/D+}(x)$ and $f_{LT}^{\,s/D+}(x)$ for the deuteron at $Q^{2}=2.5$ GeV2 by using the twist-2 Wandzura-Wilczek-like relation. This is the first numerical estimate of the tensor-polarized twist-3 distributions. Their functional forms and magnitudes are roughly similar to the twist-2 distributions $f_{1LL}^{\,u/D+}(x)=f_{1LL}^{\,d/D+}(x)$ and $f_{1LL}^{\,s/D+}(x)$. We also confirmed that the Burkhardt-Cottingham-like sum rule is satisfied numerically. The twist-3 distributions $f_{LT}^{\,q/D+}(x)$ should be investigated for the deuteron by the future semi-inclusive DIS and Drell-Yan measurements.
@@ Acknowledgments
The authors were partially supported by a research program of the Chinese Academy of Sciences.

### v2

@@ 1 Introduction
The JLab experiments for the tensor-polarized deuteron will be done in the near future. $Q^{2}(=-q^{2})$ values, where $q$ is the momentum transfer in the electron scattering, are not much larger than the typical hadronic energy scale of the order of 1 GeV2 in the JLab measurements. This fact suggests that higher-twist terms should be understood in extracting even the twist-2 tensor-polarized structure functions. Now, the theoretical formalism is completed to calculate cross sections of inclusive and semi-inclusive DIS (SIDIS) processes for spin-1 hadrons [25], and these processes could be used for measuring the tensor-polarized twist-3 distribution functions $f_{LT}(x)$. In addition, Drell-Yan processes could be used for studying $f_{LT}(x)$ at the hadron accelerator facilities [26]. Considering this situation, we show possible twist-3 distribution functions $f_{LT}(x)$ theoretically for the deuteron in this work. Because the distributions $f_{LT}(x)$ are written in terms of the twist-2 distributions $f_{1LL}(x)$ by the WW-like relation and a parametrization exists for $f_{1LL}(x)$, actually tensor-polarized PDFs $\delta_{T}q(x)$ [27], we can calculate $f_{LT}(x)$ by using this twist-2 relation. The obtained $f_{LT}(x)$ should satisfy the BC-like sum rule.
@@ 2 Twist-2 and 3 tensor-polarized parton distribution functions of the spin-1 deuteron
Here, $k$, $P$, and $T$ are the quark momentum, the deuteron momentum, and the tensor polarization of the deuteron, $\psi$ is the quark field, and $W^{[c]}$ is the gauge link with the integral path $c$ necessary for the color gauge invariance. The tensor $T^{\mu\nu}$ is symmetric and traceless, and it satisfies $P_{\mu}T^{\mu\nu}=0$. Then, integrating the function over the momenta $k^{-}$ and the transverse momentum $\vec{k}_{T}$ and fixing the $k^{+}$ component, we obtain the collinear correlation function
@@
For the spin-1 deuteron, the spin tensor $T^{\mu\nu}$ is parametrized as [18, 19]
$\hskip-5.69046ptT^{\mu\nu}=\frac{1}{2}\left[\frac{4}{3}S_{LL}\frac{(P\cdot n)^{2}}{M_{D}^{2}}\bar{n}^{\mu}\bar{n}^{\nu}-\frac{2}{3}S_{LL}(\bar{n}^{\{\mu}n^{\nu\}}-g_{T}^{\mu\nu})\right$.
$\hskip-9.95863pt\left.+\frac{1}{3}S_{LL}\frac{M_{D}^{2}}{(P\cdot n)^{2}}n^{\mu}n^{\nu}+\frac{P\cdot n}{M_{D}}\bar{n}^{\{\mu}S_{LT}^{\nu\}}-\frac{M_{D}}{2P\cdot n}n^{\{\mu}S_{LT}^{\nu\}}+S_{TT}^{\mu\nu}\right],$ (3)
where $M_{D}$ is the deuteron mass and $a^{\{\mu}b^{\nu\}}$ is defined by $a^{\{\mu}b^{\nu\}}=a^{\mu}b^{\nu}+a^{\nu}b^{\mu}$. The lightcone vectors $n$ and $\bar{n}$ are defined by
[…]
The $g_{T}^{\mu\nu}$ is the transverse projector, $g_{T}^{\mu\nu}=g^{\mu\nu}-\bar{n}^{\{\mu}n^{\nu\}}$. The tensor polarization along the longitudinal axis is expressed by $S_{LL}$. The parameters $S_{LT}^{\mu}$ and $S_{TT}^{\mu\nu}$ indicate polarization differences along the axes between the longitudinal and transverse directions and along the transverse axes, respectively [18, 19]. They satisfy the relations $S_{LT}\cdot\bar{n}=S_{LT}\cdot n=0$, $S_{TT}^{\mu\nu}\bar{n}_{\mu}=S_{TT}^{\mu\nu}n_{\mu}=0$ and ${g_{T}}_{\mu\nu}S_{TT}^{\mu\nu}=0$.
@@
$\Phi(x_{D})=$ $\frac{1}{2}\,\bigg[\,S_{LL}\,\not{\bar{n}}\,f_{1LL}(x_{D})+\frac{M_{D}}{P\cdot n}\,S_{LL}\,e_{LL}(x_{D})$
@@
In the polarized structure functions of the nucleon, the WW relation has been often used for estimating the twist-3 structure function $g_{2}$ in terms of the twist-2 one $g_{1}$. In addition, the BC sum rule is also useful for the estimate of $g_{2}$. For a review of the original WW relation and the BC sum rule in the spin-1/2 case, one may look at Ref. [28]. In a similar way to the derivations of these relations, the WW-like relation was obtained for the distribution functions $f_{LT}(x_{D})$ and $f_{1LL}(x_{D})$ by neglecting dynamical twist-3 terms as  [20]
@@
and the BC-like sum rule was derived
$\int_{0}^{1}dx_{D}\,f_{2LT}^{\,q+}(x_{D})=0,\ \ \ f_{2LT}^{\,q+}(x)\equiv\frac{2}{3}f_{LT}^{\,q+}(x_{D})-f_{1LL}^{\,q+}(x_{D}).$ (7)
Here, the quark flavor type is shown by $q$ in the superscript as $f^{\,q+}$, which is defined by
@@
The original WW relation was derived in terms of the operator product expansion (OPE) with local operators [28, 29]. In our case, the WW-like relation was obtained by nonlocal operators with the tensor polarization parameters [20]. It should be noted that the same WW-like relation is obtained also by the local OPE with the mixed-symmetry operator $\bar{\psi}(0)\gamma^{\,[\sigma}D^{\,\{\mu_{1}]}\cdots D^{\,\mu_{n-1}\}}\psi(0)$ and the symmetry one $\bar{\psi}(0)\gamma^{\,\{\sigma}D^{\,\mu_{1}}\cdots D^{\,\mu_{n-1}\}}\psi(0)$, where $\cdots$] indicates the antisymmetric combination. Their matrix elements are decomposed into the mixed-symmetric and symmetric terms in terms of the momentum $P^{\,\mu}$ and the tensor $T^{\mu\nu}$. Then, the WW-like relation is obtained in a similar way as explained in Sec. 2 of Ref. [20]. Therefore, the current WW-like relation is consistent with the local OPE derivation.
@@ 3.2 Twist-3 distribution functions $f_{LT}(x)$ by using the twist-2 relations
Using the PDFs in Fig. 1, we calculated the tensor-polarized twist-2 distribution functions $f_{1LL}^{\,q/D+}$, which are expressed by the PDFs $\delta_{T}q^{D}$ and $\delta_{T}\bar{q}^{D}$ as [1, 20]
@@
In order to confirm the BC-like sum rule in Eq. (17), the distributions $f_{1LL}^{\,q/D+}(x)$ and $f_{LT}^{\,q/D+}(x)$ are combined to obtain to $f_{2LT}^{\,q/D+}(x)$ at $Q^{2}=2.5$ GeV2 in Fig. 4. According to the sum rule, the integrals of these distributions should vanish. In fact, the positive and negative areas in this figure cancel with each other. However, finite distributions still exist at $x\leq 10^{-3}$ in Fig. 4 and the integrals become much smaller, of the order of $10^{-5}$, if the integrals are calculated from $x_{\min}=10^{-9}$.
The higher-twist distributions are generally suppressed by $1/Q$ in cross sections in comparison with the leading-twist ones. However, the $Q^{2}$ values of the JLab experiment are not very large, so that there is a possibility that the higher-twist distributions $f_{LT}^{\,q/D+}(x)$ could be measured experimentally by the inclusive and semi-inclusive DIS (SIDIS) [25] in addition to the leading-twist ones $f_{1LL}^{\,q/D+}(x)$ [3]. The SIDIS cross section is expressed by various functions denoted as [25]
[…]
The twist-3 distribution function $f_{LT}^{\,q/D+}(x)$ is contained in the function $F_{U(LT)}^{\,\cos\phi_{LT}}$. Namely, using the unpolarized charged-lepton beam and the target with the tensor polarization (LT) and looking at the dependence of $\cos\phi_{LT}$ in the cross section, we can measure the function $f_{LT}^{\,q/D+}(x)$. Here, $\phi_{LT}$ is the azimuthal angle of tensor polarization $S_{LT}^{\mu}$. According to Eqs. (4.111) and (4.113)-(4.116) in Ref. [25], $f_{LT}^{\,q/D+}(x)$ is contained in the structure functions $b_{2\text{--}4}$ in inclusive DIS. Thus, the distribution $f_{LT}^{\,q/D+}(x)$ could be measured at JLab [3, 4] and future EICs [12, 13].
Furthermore, Drell-Yan processes could also be used for finding the twist-3 distributions $f_{LT}^{\,q/D+}(x)$ [26]. In the proton-deuteron Drell-Yan process with the tensor-polarized deuteron, the function $f_{LT}^{\,q/D+}(x)$ appears in the cross section with the angular dependence of $\cos\hat{\phi}$, where $\hat{\phi}$ is $\hat{\phi}=\phi-\phi_{s}$ with the azimuthal angle $\phi$ of $\mu^{-}$ and the transverse-vector angle $\phi_{s}$ of $S_{LT}^{\mu}$. Therefore, there is a possibility that $f_{LT}^{\,q/D+}(x)$ could be measured in the Fermilab experiment [8, 9] and in other experiments at the hadron accelerator facilities [10, 11].
@@ 4 Summary
Using the parametrization for the tensor-polarized PDFs $\delta_{T}q^{D}(x)$ and $\delta_{T}\bar{q}^{D}(x)$, we calculated the tensor-polarized twist-3 PDFs $f_{LT}^{\,u/D+}(x)=f_{LT}^{\,d/D+}(x)$ and $f_{LT}^{\,s/D+}(x)$ for the deuteron at $Q^{2}=2.5$ GeV2 by using the twist-2 Wandzura-Wilczek-like relation. This is the first numerical estimate of the tensor-polarized twist-3 distributions. Their functional forms and magnitudes are roughly similar to the twist-2 distributions $f_{1LL}^{\,u/D+}(x)=f_{1LL}^{\,d/D+}(x)$ and $f_{1LL}^{\,s/D+}(x)$. We also confirmed that the Burkhardt-Cottingham-like sum rule is satisfied numerically. The twist-3 distributions $f_{LT}^{\,q/D+}(x)$ should be investigated for the deuteron by the future inclusive DIS, semi-inclusive DIS, and Drell-Yan measurements.
@@ Acknowledgments
The authors were partially supported by a research program of the Chinese Academy of Sciences. KK was also supported by the Gansu-province postdoctoral foundation. They thank Qin-Tao Song for suggestions.